# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v2

@@ 5 Experiments
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution (8×8$\times$8 ×)
@@ 5.1 Experimental Setting
Compared methods. We compare LD-SMC with several recent SoTA baseline methods. All methods were evaluated in a similar experimental setup using the same latent diffusion model to ensure fairness in the comparisons. (1) DPS (Chung et al., 2023b), which introduces correction to the sampling process of the diffusion through the posterior mean estimator; (2) TDS (Wu et al., 2024), which uses the twisting technique for approximate sequential Monte Carlo sampling; since DPS and TDS were designed for pixel-space diffusion models, we term the adapted versions of them to the latent space LDPS and LTDS, respectively. (3) ReSample (Song et al., 2024), which applies an optimization procedure during the sampling process to match the approximate posterior mean to the label followed by a resampling step; (4) PSLD (Rout et al., 2023), which introduces a correction term to the DPS step to “glue” ${\mathbf{z}}_{0}$; (5) PFLD (Nazemi et al., 2024), a particle filtering approach for LDM based on PSLD update for the proposal distribution; (6) LatentDAPS which decouples the latents in the sampling trajectory. In all experiments, we evaluated LD-SMC with $N=\{1,5\}$ particles. Using one particle incurs a computational cost similar to that of LDPS (see Table 4 in the appendix). LTDS and PFLD were evaluated with $N=5$ particles. All methods were tested on the following inverse problem tasks: box inpainting, Gaussian deblurring, super-resolution (8×8$\times$8 ×) and free-form inpainting. We followed the setup in (Chung et al., 2023b) for the first three tasks and the setup in (Saharia et al., 2022) for the last one with a few changes to the default hyperparameters. In inpainting tasks, the masks were randomly sampled for each new sample.
@@
(b) Effect of $N$
Figure 5: FID and PSNR values when varying $s$ (left) and the number of particles $N$ (right) on ImageNet box (left) and free-form (right) inpainting tasks.
@@ 6 Conclusion
In this study, we presented LD-SMC, a novel method for solving inverse problems in the latent space of diffusion models using SMC. LD-SMC augment the model with auxiliary observations, one per each timestep, and use these observations to guide the sampling process as part of the backward diffusion process. This framework can be seen as applying one step of blocked Gibbs sampling. To perform SMC sampling, we suggested a novel weighing scheme and proposal distributions. Both are based on information from the auxiliary labels and the true label ${\mathbf{y}}_{0}$. Empirically, we validated LD-SMC against strong baseline methods on common benchmarks. The results suggest that LD-SMC can improve the performance over baseline methods, especially in cases where extrapolation is needed (e.g., in inpainting tasks). one limitation of our approach is related to computational demand. The sampling time and the memory demand increase with the number of particles and Gibbs iterations. This effect can be partially mitigated by taking fewer particles and GPU parallelization.
@@ Appendix C Asymptotic Accuracy of LD-SMC
Here we provide a proof that ${\mathbf{z}}_{0}$ samples can be made arbitrarily accurate according to our model design. Our proof is composed of three parts, all of which have been previously established in the literature. We will restate them here and accommodate them to our setting. The three parts are, (1) augmenting the model with auxiliary random variables, ergo performing a completion of the desired marginal density $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ (Definition 10.3 in (Robert & Casella, 1999)), (2) Proving asymptotic accuracy of the SMC procedure following a similar lines of Theorem 2 in (Wu et al., 2024), (3) Showing sufficient conditions that the Markov chain generated by the Gibbs sampling procedure is ergodic and hence $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ is the limiting distribution from which ${\mathbf{z}}_{0}$’s are sampled (Theorem 10.6 in (Robert & Casella, 1999)).
@@
[…]
@@
To prove the asymptotic accuracy of the SMC procedure in LD-SMC, we adopt the formulation of (Wu et al., 2024). In what follows, to prevent cluttered notation, we drop the index of the Gibbs iteration and note that all quantities are those of a specific iteration. Specifically, here we first reiterate the following $3$ important quantities, the prior distribution $p_{\theta}({\mathbf{z}}_{0:T})$, the proposal distributions $\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})$, and the weighting functions $w_{t}({\mathbf{z}}_{t},{\mathbf{z}}_{t+1})$ and then proceed to the proof.
@@
Now that we established necessary conditions to sample from $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$, the following theorem from (Chopin et al., 2020) characterizes the conditions under which SMC algorithms converge. We adopt the formulation of (Wu et al., 2024) and present it here, adapted to our case.
@@
[…]
@@
[…]
Note that here, unlike (Wu et al., 2024), we do not need to assume that $\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})=p_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})$ due to our revised posterior distribution. Here, however, we add assumption (c). Assumptions (b) and (c) are the strongest assumptions as they may not apply even for linear transformations. But, for sufficiently smooth decoder and diffusion model in the input (${\mathbf{z}}_{t}$ or $\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)$) with uniformly bounded gradients, the assumptions will hold.
@@
$\log~{}\frac{p_{\theta}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}{\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})}$ $=\log~{}\frac{|2\pi\sigma_{t+1}^{2}{\mathbf{I}}|^{-0.5}\exp\{-(2\sigma_{t+1}^{2})^{-1}||{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t}||_{2}^{2}\}}{|2\pi\tilde{\sigma}_{t+1}^{2}{\mathbf{I}}|^{-0.5}\exp\{-(2\tilde{\sigma}_{t+1}^{2})^{-1}||{\mathbf{z}}_{t}-{\mathbf{m}}_{t}||_{2}^{2}\}}$
=$\stackrel{{\scriptstyle\mathclap{\mbox{c}}}}{{=}}-\frac{1}{2}[\sigma_{t+1}^{-2}||{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t}||_{2}^{2}-\tilde{\sigma}_{t+1}^{-2}||{\mathbf{z}}_{t}-{\mathbf{m}}_{t}||_{2}^{2}]$
@@
$=-\frac{1}{2}[(\sigma_{t+1}^{-2}-\tilde{\sigma}_{t+1}^{-2})||{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t}||_{2}^{2}-2\tilde{\sigma}_{t+1}^{-2}\langle{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t},\bm{\mathbf{\mu}}_{t}-{\mathbf{m}}_{t}\rangle-\tilde{\sigma}_{t+1}^{-2}||\bm{\mathbf{\mu}}_{t}-{\mathbf{m}}_{t}||_{2}^{2}]$
@@
=$\stackrel{{\scriptstyle\mathclap{\mbox{c}}}}{{=}}-\frac{1}{2}[(\sigma_{t+1}^{-2}-\tilde{\sigma}_{t+1}^{-2})||{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t}||_{2}^{2}-2\tilde{\sigma}_{t+1}^{-2}\langle{\mathbf{z}}_{t}-\bm{\mathbf{\mu}}_{t},\bm{\mathbf{\mu}}_{t}-{\mathbf{m}}_{t}\rangle]$
@@
$\leq\frac{\tilde{\sigma}_{t+1}^{-4}||\bm{\mathbf{\mu}}_{t}-{\mathbf{m}}_{t}||_{2}^{2}}{2(\sigma_{t+1}^{-2}-\tilde{\sigma}_{t+1}^{-2})}<\infty.$
Where the last inequality is again due to assumptions (b) and (c). Here, =$\stackrel{{\scriptstyle\mathclap{\mbox{c}}}}{{=}}$ denotes an equality up to a constant. ∎
@@ Appendix D Proposal Distribution Scaling Coefficients
Let ${\mathbf{g}}_{t}^{1}\coloneqq\nabla_{{\mathbf{z}}_{t}}||{\mathbf{y}}_{0}-{\mathcal{A}}({\mathcal{D}}(\bar{{\mathbf{z}}}_{0}({\mathbf{z}}_{t})))||_{2}^{2}$, and ${\mathbf{g}}_{t}^{2}\coloneqq\nabla_{\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)}||{\mathbf{y}}_{t}-{\mathcal{A}}({\mathcal{D}}(\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)))||_{2}^{2}$. We set the scaling coefficients $\gamma_{t}$ and $\lambda_{t}$ according to the following scheme: $\gamma_{t}=\mathbbm{1}_{[t\geq s]}\cdot\kappa_{1}\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}+\mathbbm{1}_{[t<s]}\cdot\kappa_{2}\cdot(1-\rho)\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{1}||_{2}^{2},1)}$, and $\lambda_{t}=\mathbbm{1}_{[t<s]}\cdot\kappa_{2}\cdot\rho\cdot\frac{1}{max(||{\mathbf{g}}_{t}^{2}||_{2}^{2},1)}$, where $s,\rho,\kappa_{1},\kappa_{2}$ are all hyper-parameters.
@@ Appendix H Full Results
Table 9: ImageNet. Super-resolution (8×8$\times$8 ×) on $1024$ test examples.
@@
Table 13: FFHQ. Super-resolution (8×8$\times$8 ×) on $1024$ test examples.

### v3

@@ 5 Experiments
Inpainting (Box) Inpainting (Free-Form) Gaussian Deblurring Super-Resolution ($8\times$)
@@ 5.1 Experimental Setting
Compared methods. We compare LD-SMC with several recent SoTA baseline methods. All methods were evaluated in a similar experimental setup using the same latent diffusion model to ensure fairness in the comparisons. (1) DPS (Chung et al., 2023b), which introduces correction to the sampling process of the diffusion through the posterior mean estimator; (2) TDS (Wu et al., 2023), which uses the twisting technique for approximate sequential Monte Carlo sampling; since DPS and TDS were designed for pixel-space diffusion models, we term the adapted versions of them to the latent space LDPS and LTDS, respectively. (3) ReSample (Song et al., 2024), which applies an optimization procedure during the sampling process to match the approximate posterior mean to the label followed by a resampling step; (4) PSLD (Rout et al., 2023), which introduces a correction term to the DPS step to “glue” ${\mathbf{z}}_{0}$; (5) PFLD (Nazemi et al., 2024), a particle filtering approach for LDM based on PSLD update for the proposal distribution; (6) LatentDAPS which decouples the latents in the sampling trajectory. In all experiments, we evaluated LD-SMC with $N=\{1,5\}$ particles. Using one particle incurs a computational cost similar to that of LDPS (see Table 4 in the appendix). LTDS and PFLD were evaluated with $N=5$ particles. All methods were tested on the following inverse problem tasks: box inpainting, Gaussian deblurring, super-resolution ($8\times$) and free-form inpainting. We followed the setup in (Chung et al., 2023b) for the first three tasks and the setup in (Saharia et al., 2022) for the last one with a few changes to the default hyperparameters. In inpainting tasks, the masks were randomly sampled for each new sample.
@@
(b) Effect of NN
Figure 5: FID and PSNR values when varying $s$ (left) and the number of particles NN (right) on ImageNet box (left) and free-form (right) inpainting tasks.
@@ 6 Conclusion
In this study, we presented LD-SMC, a novel method for solving inverse problems in the latent space of diffusion models using SMC. LD-SMC augment the model with auxiliary observations, one per each timestep, and use these observations to guide the sampling process as part of the reverse diffusion process. This framework can be seen as applying one step of blocked Gibbs sampling. To perform SMC sampling, we suggested a novel weighing scheme and proposal distributions. Both are based on information from the auxiliary labels and the true label ${\mathbf{y}}_{0}$. Empirically, we validated LD-SMC against strong baseline methods on common benchmarks. The results suggest that LD-SMC can improve the performance over baseline methods, especially in cases where extrapolation is needed (e.g., in inpainting tasks). one limitation of our approach is related to computational demand. The sampling time and the memory demand increase with the number of particles and Gibbs iterations. This effect can be partially mitigated by taking fewer particles and GPU parallelization.
@@ Appendix C Asymptotic Accuracy of LD-SMC
Here we provide a proof that ${\mathbf{z}}_{0}$ samples can be made arbitrarily accurate according to our model design. Our proof is composed of three parts, all of which have been previously established in the literature. We will restate them here and accommodate them to our setting. The three parts are, (1) augmenting the model with auxiliary random variables, ergo performing a completion of the desired marginal density $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ (Definition 10.3 in (Robert & Casella, 1999)), (2) Proving asymptotic accuracy of the SMC procedure following a similar lines of Theorem 2 in (Wu et al., 2023), (3) Showing sufficient conditions that the Markov chain generated by the Gibbs sampling procedure is ergodic and hence $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$ is the limiting distribution from which ${\mathbf{z}}_{0}$’s are sampled (Theorem 10.6 in (Robert & Casella, 1999)).
@@
(Robert & Casella, 1999). Given a probability density $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$, a density $g$ that satisfies $\int g({\mathbf{z}}_{0},{\mathbf{z}}_{1:T},{\mathbf{y}}_{1:T}|{\mathbf{y}}_{0})d{\mathbf{z}}_{1:T}d{\mathbf{y}}_{1:T}$ is called a completion of $p_{\theta}({\mathbf{z}}_{0}|{\mathbf{y}}_{0})$.
@@
To prove the asymptotic accuracy of the SMC procedure in LD-SMC, we adopt the formulation of (Wu et al., 2023). In what follows, to prevent cluttered notation, we drop the index of the Gibbs iteration and note that all quantities are those of a specific iteration. Specifically, here we first reiterate the following $3$ important quantities, the prior distribution $p_{\theta}({\mathbf{z}}_{0:T})$, the proposal distributions $\pi_{t}({\mathbf{z}}_{t}|{\mathbf{z}}_{t+1})$, and the weighting functions $w_{t}({\mathbf{z}}_{t},{\mathbf{z}}_{t+1})$ and then proceed to the proof.
@@
Now that we established necessary conditions to sample from $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$, the following theorem from (Chopin et al., 2020) characterizes the conditions under which SMC algorithms converge. We adopt the formulation of (Wu et al., 2023) and present it here, adapted to our case.
@@
$\lim_{N\rightarrow\infty}\sum_{i=1}^{N}w_{0}^{(i)}\phi({\mathbf{z}}_{0:T}^{(i)})=\int\phi({\mathbf{z}}_{0:T})\nu_{0}({\mathbf{z}}_{0:T})d{\mathbf{z}}_{0:T}$
@@
Then ${\mathbb{P}}_{N}$ converges setwise to $p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})$ with probability one, that is for every set $E$, $\lim_{N\rightarrow\infty}{\mathbb{P}}_{N}(E)=\int_{E}p_{\theta}({\mathbf{z}}_{0:T}|{\mathbf{y}}_{0:T})d{\mathbf{z}}_{0:T}$.
Note that here, unlike (Wu et al., 2023), we do not need to assume that $\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})=p_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{0})$ due to our revised posterior distribution. Here, however, we add assumption (c). Assumptions (b) and (c) are the strongest assumptions as they may not apply even for linear transformations. But, for sufficiently smooth decoder and diffusion model in the input (${\mathbf{z}}_{t}$ or $\bm{\mathbf{\mu}}_{\theta}({\mathbf{z}}_{t+1},t+1)$) with uniformly bounded gradients, the assumptions will hold.
@@
[…]
$\text{Notice, }||\bm{\mathbf{\mu}}_{t}-{\mathbf{m}}_{t}||_{2}^{2}=||\gamma_{t}\nabla_{{\mathbf{z}}_{t+1}}\log~\bar{p}_{\theta}({\mathbf{y}}_{0}|{\mathbf{z}}_{t+1})+\lambda_{t}\nabla_{\bm{\mathbf{\mu}}_{t}}\log~q_{\theta}({\mathbf{y}}_{t}|{\mathbf{z}}_{t+1})||_{2}^{2}<\infty$
@@
[…]
@@
$\text{Apply the inequality }-\frac{a}{2}x^{2}+bx\leq\frac{b^{2}}{2a}\text{for }a=\sigma_{t+1}^{-2}-\tilde{\sigma}_{t+1}^{-2}>0\text{ by assumption (d)}$
@@
Where the last inequality is again due to assumptions (b) and (c). Here, $\stackrel{{\scriptstyle\mathclap{\mbox{c}}}}{{=}}$ denotes an equality up to a constant. ∎
@@ Appendix D Proposal Distribution Scaling Coefficients
[…]
@@ Appendix H Full Results
Table 9: ImageNet. Super-resolution ($8\times$) on $1024$ test examples.
@@
Table 13: FFHQ. Super-resolution ($8\times$) on $1024$ test examples.